Suppose 8 bit registers have following contents
X=00001111
Y=10101010
Z= 11011011
W=00110011
What will be the 8 bit values of each register after execution of following sequences of
microoperations ?
X ←
1
Expert's answer
2021-09-25T05:20:43-0400

The given 8 digit bits are

X=00001111

Y=10101010

Z= 11011011

W = 00110011


X=00001111 = 15

Y=10101010 = 170

z = 11011011 = 219

w = 00110011 = 51


x + y = 10111001

z+w = 00001110

Now x+y+z+w = 11000111